A company analyses its profits for the three consecutive years, from 2017-2019. The profit made in 2017 is three times that made in 2018 and the profit made in 2019 is R5 million rand more than the profit made in 2012. The total profit for the three years is 80 million. Calculate the profit in each year
1
Expert's answer
2020-06-15T15:28:39-0400

A company analyses its profits for the three consecutive years, from 2017-2019. The profit made in 2017 is three times that made in 2018 and the profit made in 2019 is R5 million rand more than the profit made in 2018. The total profit for the three years is 80 million. Calculate the profit in each year

Let "x=" the profit made in 2018. Then the profit made in 2017 is "3x," and the profit made in 2019 is "x+5." The total profit for the three years is 80 million


"3x+x+x+5=80""5x=80-5""5x=75""x=15"

The profit made in 2017 is R45 million rand.

The profit made in 2018 is R15 million rand.

The profit made in 2019 is R20 million rand.